OpenStudy (anonymous):

Solve. Express your answer in simplest form. 3 3/5÷ 3/5 A.25/ 57 B. 3/19 C.6 1/3 D.2 7/25

OpenStudy (studygurl14):

First step, convert \(3\large\frac{3}{5}\) into an improper fraction. Also remember that when dividing a fraction, you multiply by the reciprocal. Example: \(\large\frac{1}{2}\div\frac{3}{4}=\frac{1}{2}\times\frac{4}{3}\)